OpenStudy (dancing_stars23):

1) The average kinetic energy of 1 mole of a gas at standard temperature is: 0 J 3.40 x 103 J 1.51 x 103 J 3.36 x 101 J 2) How does the standard atmospheric pressure exhibited at sea level compare to the atmospheric pressure exhibited at the top of a tall mountain? It is higher than the atmospheric pressure at the mountain elevation. It is the same as the atmospheric pressure at the mountain elevation. It is lower than the atmospheric pressure at the mountain elevation. This cannot be determined from the information given.

OpenStudy (academicgurusinc):

1) Kinetic Energy = (3/2)*n*R*T Here n = 1 mole, R = 8.134 J/mol K ,T = 273.16° K Substituting, K.E. = (3/2)*1*8.314*273.16 = 3406.57836 J = 3.40x10^3 2) It is higher than the atmospheric pressure at the mountain elevation. Think of gravity pushing down on the world. Air has a small amount of weight to it. However, if you were to stack several layers of air on top of one another, the bottom would exhibit a heavier weight than the top layer. An analogy would be a stack of pancakes. The top pancake experiences less pressure in comparison to the bottom pancake with several pancakes piled on top of it.

OpenStudy (academicgurusinc):

I hope that helps. If you found this helpful, I encourage you to subscribe to our youtube channel ( https://www.youtube.com/channel/UCYiI7SmkU4_vhdSzKBWsifg), to stay current with all of our new videos. Regards, Academic Gurus Inc. Twitter (Academic_Gurus) Facebook (AcademicGurusInc) Youtube (Academic Gurus Inc)